Suppose the following transactions occur during the current year:

1. Musashi orders 50 bottles of wine from a French distributor at a price of $30 per bottle.
2. A U.S. company sells 200 textbooks to a Canadian company at $45.00 per textbook.
3. Sean, a U.S. citizen, pays $1,500 for a laptop he orders from Microell (a U.S. company).

Complete the following table by indicating how the combined effects of these transactions will be reflected in the U.S. national accounts for the current year.

Hint: Be sure to enter a “0” if none of the transactions listed are included in a given category and to enter a minus sign when the balance is negative.

Amount

(Dollars)

Consumption
Investment
Government Purchases
Imports
Exports
Net Exports
Gross Domestic Product (GDP)
Answer #1

Consumption = Sean purchase of the laptop from the US company = 1500

Investment = 0

Government purchases = 0

Imports = Mushashi order = 50*30 = 1500

Exports = Textbook sales = 200*45 = 9000

Net Exports = Exports - Imports = 9000-1500 = 7500

Gross domestic product(GDP) = C+I+G+NX = 1500+0+0+7500 = 9000